OpenStudy (anonymous):

Can someone please factor this trinomial for me?? Thank you!!] Please show ur work on how you got the answer. 2a2+19a+9

hero (hero):

2a^2+19a+9 =2a^2+18a+1a+9 =2a(a+9)+1(a+9) =(a+9)(2a+1)